estion lordu z #20: Module 10 Problems Question 126 PM AIA Inc. is looking to manage its cash position using the EOQ model. The company is consuming cash at the rate of $6300 per day, and is open for business 365 ays n the year Each tme the m sells securites to obtain the cash it costs them $210. The interest rate is 2.90% what are the annual order costs associated with the EOQ? CHECK ANSWER
Answer #1
EOQ = square root of (2* annual consumption*transaction cost/Interest rate)
Annual consumption =6300*365 2299500
Transaction cost 210
Interest rate 2.90%
EOQ = =square root of (2*2299500*210/2.9%)
EOQ = 182491
Average cash balance = 182491/2       91,245.69
Number of transaction = 2299500/182491              12.60
Transaction cost = 12.6*210         2,646.12